Is (2,3) the solution to the linear system defined by
y=3x-3
2x+4y=16

Answer by bam878s(77)

All you need to do is plug x = 2 and y = 3 into the equations and see if they are true.
3(2) - 3 = 6 - 3 = 3 this equation check out
2(2) + 4(3) = 4 + 12 = 16 this equation also checks out
so (2,3) is a solution